OpenStudy (anonymous):

Derivatives: y=x^1/8(6^x)

OpenStudy (anonymous):

Product rule! f = x^(1/8) and g = 6^x. Now find f' and g' and plug it into the formula: f'g + g'f

OpenStudy (anonymous):

@BluFoot ok... I got f' = (x/8)^-7/8 if that makes sense... but how exactly do you work out the g'?

OpenStudy (anonymous):

Use this rule: derivative of a^x = a^x * ln(a) where a is any number

OpenStudy (anonymous):

@BluFoot I got: g' = 6^x * In ^ 6 so my final answer is: f'g + g'f = \[x/8^(- 7/8) * 6^x + 6^x * In 6 * x^ (1/8)\]

OpenStudy (anonymous):

Perfect, very good! :)

OpenStudy (anonymous):

@BluFoot thank you so much!!!!! Just did another example and got the correct answer. I think i've grasped that concept now. :)

OpenStudy (anonymous):

Great! You're very welcome! It's much easier when you make a little table or something where you write f = this, g= that, f' = derivative of this, g'= derivative of that. Then you just combine the 4 terms and you have your answer :)

OpenStudy (anonymous):

I'll be sure to keep that in mind when my test comes up. *If only my lecturer made things this simplified...life would be so much better.* Thank you once again. :)